# Pushing the Limits of Sparsity: A Bag of Tricks for Extreme Pruning

### v3

@@ Abstract
Pruning of deep neural networks has been an effective technique for reducing model size while preserving most of the performance of dense networks, crucial for deploying models on memory and power-constrained devices. While recent sparse learning methods have shown promising performance up to moderate sparsity levels such as 95% and 98%, accuracy quickly deteriorates when pushing sparsities to extreme levels due to unique challenges such as fragile gradient flow. In this work, we explore network performance beyond the commonly studied sparsities, and propose a collection of techniques that enable the continuous learning of networks without accuracy collapse even at extreme sparsities, including 99.90%, 99.95% and 99.99% on ResNet architectures. Our approach combines 1) Dynamic ReLU phasing, where DyReLU initially allows for richer parameter exploration before being gradually replaced by standard ReLU, 2) weight sharing which reuses parameters within a residual layer while maintaining the same number of learnable parameters, and 3) cyclic sparsity, where both sparsity levels and sparsity patterns evolve dynamically throughout training to better encourage parameter exploration. We evaluate our method, which we term Extreme Adaptive Sparse Training (EAST) at extreme sparsities using ResNet-34 and ResNet-50 on CIFAR-10, CIFAR-100, and ImageNet, achieving significant performance improvements over state-of-the-art methods we compared with.
@@ 1 Introduction
Network pruning [9, 10, 16, 22, 19, 15] is a widely-used technique for reducing a network’s parameters and compressing its size. Reducing model sizes is crucial for deploying models on edge devices with limited resources. Conventionally, pruning methods have focused on reducing parameters from pre-trained models. However, it requires at least as much computation as training a dense model as it must converge before pruning takes place. The Lottery Ticket Hypothesis [6, 7, 23] gives theoretical foundation that subnetworks have the potential to reach full performance even when trained from an initially sparse state. This insight has recently gained much traction in sparse training, a paradigm in which sparse networks are trained from scratch without the need for dense pre-training.
@@
Sparse training methods can be broadly classified into two categories: static sparse training (SST), also sometimes referred to as Pruning at Initialization (PaI), where the sparsity pattern is pre-determined at initialization and remains fixed throughout training [17, 33, 30, 2], and dynamic sparse training (DST), where the sparsity pattern continuously evolves during training [25, 3, 4, 34]. These methods perform well at moderate sparsity commonly up to 98%, but experience rapid accuracy degradation beyond this due to limitations such as layer collapse and gradient issues [31, 33]. There is an increasing demand to push towards ’extreme sparsities’, particularly in resource-constrained settings where substantial computational and memory benefits may justify some accuracy trade-offs. Our work is motivated by this practical consideration, and thus we emphasize that our primary focus is on the deployable model used for inference rather than training efficiency. Despite this need, extreme sparsities remain relatively understudied, with only a few methods attempting to maintain usable accuracy at these extreme levels [31, 2, 28].
@@
Our work proposes a collection of adaptive methods that when combined can achieve meaningful performance at extreme sparsities. By leveraging three core strategies - 1) phased Dynamic ReLU activation, 2) weight sharing, 3) cyclic sparsity scheduling - we present a flexible framework that is capable of optimizing performance at extreme sparsities. Each component independently enhances the sparse model’s capacity as seen in our ablation studies, where individual strategy improves performance on their own. However, when used together, these methods can cohesively achieve an even stronger performance. This framework, which we term Extreme Adaptive Sparse Training (EAST), offers a modular approach that can be adapted to existing DST frameworks, effectively postponing their total performance collapse, as shown in Figure 2. Through empirical analysis, we find that EAST prevents gradient vanishing and encourages parameter exploration, achieving substantial and meaningful performance gain over our baselines. Our contribution and key achievements can be summarized as follows:
@@
We perform extensive empirical evaluation of EAST on image classification tasks, and demonstrate that it can achieve superior performance at sparsities beyond 99.90%, surpassing all previous methods tested.
We explore and demonstrate that DyReLU when used as a drop-in replacement for ReLU at initialization at then replaced completely can still improve the parameter exploration at extreme sparsities, leading to higher performance than sparse models trained using ReLU only.
To the best of our knowledge, EAST is the first DST method to dynamically adjust both sparsity patterns and sparsity levels throughout training (as opposed to just the patterns). Our cyclic sparsity schedule allows models to explore parameters and stabilize at extreme sparsity, as evidenced by our ablation studies.
@@ 2 Related Work
Dynamic Sparse Training. Early work by Mocanu et al. demonstrated the viability of DST through SET [25], which used magnitude pruning and random weight regrowth. Subsequent criteria for regrowing connections have been developed, such as using the momentum of parameters in SNFS [3] or their gradient in RigL [4] to determine the salience of deactivated connections. GraNet [20] adapted RigL’s approach by starting with half the parameters and gradually increasing to target sparsity instead of starting from the target sparsity. ITOP [21] provides a deeper understanding of pre-existing DST methods and enhances parameter exploration during sparse training. Supticket [34] further enhances the performance of DST by leveraging weight averaging at the late training stage. NeurRev [18] addressed the issue of dormant neurons by removing harmful negative weights. Recent advances include Top-KAST [13], which maintains an auxiliary set of weights for enhanced exploration, and BiDST [14] which formulates DST as a bi-level optimization problem for joint weight-mask optimization. MEST [36] introduces a memory-efficient sparse training framework for edge devices, while Chase [35] adapts dynamic unstructured sparsity into hardware-friendly channel-wise sparsity, both optimizing performance on resource-constrained systems without sacrificing accuracy.
@@
Extreme Sparsity Methods. Most current methods focus on sparsity levels at which pruned networks match or closely approximate the performance of their dense counterparts, a concept referred to as matching sparsity [8]. However, there is growing interest in extreme sparsities, where accuracy trade-offs become inevitable but potentially worthwhile for substantial computational gains. Yet there remains challenges, such as maintaining stable gradient flow and preventing layer collapse, which conventional pruning methods struggle to address. GraSP [33] and SynFlow [31] specifically addressed the gradient preservation and layer collapse issues at high sparsities via careful network initialization. SNIP [17] and SNIP-it [32] proposed PaI methods that showed promising results at higher sparsities, though primarily focused on static masks. De Jorge et al. [2] demonstrated that methods like SNIP and GraSP can perform worse than random pruning at 99% sparsity and beyond, and proposed iterative pruning approaches (Iter-SNIP and FORCE) that maintained meaningful performance up to 99.5% sparsity through gradual parameter elimination. DCTpS [28] introduced ”DCT plus Sparse” layers that combined a fixed dense DCT offset matrix with trainable sparse parameters to maintain propagation at extreme sparsities up to 99.99%.
@@
Activation Functions. Activation functions in sparse networks have traditionally relied on static activation functions such as ReLU [27], which applies a fixed threshold to the input. Parametric variants such as PReLU [11] introduced learnable parameters to adapt the negative slope, offering slight improvements in gradient flow. Swish [29] and Mish [24] aim to smooth gradients for better optimization but remain static throughout training. Dynamic ReLU (DyReLU) [1] introduced an adaptive activation that dynamically adjusts the slopes of a ReLU based on the input.
@@ 3.1 Activation Function
In this work, when initializing a network, we replace the standard ReLU activation with a non-ReLU activation. We tested several activations and found that DyReLU worked the best, specifically the DyReLU-B variant, which dynamically adapts and adjusts channel-wise activation coefficients during training, allowing for richer expressivity during early training. As training progresses, we gradually transition from DyReLU to standard ReLU, eliminating the additional parameters. Preliminary details about DST, DyReLU and details on phasing out DyReLU can be found in the Supplementary Materials.
Understanding Gradient Flow Through Neuron Activity. With ReLU, non-positive pre-activations would result in zero gradient and prevent the learning signal with ReLU. To better understand why DyReLU phasing improves performance at extreme sparsity, we analyzed pre-activation values throughout training. As shown in Figure 3, the model trained solely with ReLU maintains around 1.3% positive pre-activations throughout training, with many deeper layers showing none, interrupting gradient propagation.
@@ 3.2 Weight Sharing
Much research in sparse training has been about mask optimization - determining the best subset of parameters to keep. However, at extreme sparsities, gradient flow can collapase even with an optimal parameter subset. Therefore, we ask a different question: Can we increase the number of parameter paths in the computational graph while maintaining the same number of learnable parameters?
Our weight sharing mechanism strategically reuses parameters within each layer’s residual blocks, allowing the remaining active parameters to participate in multiple paths of the computational graph. If at a certain sparsity, we have remaining parameters in the original model, we keep exactly this many unique parameters to learn, but some of these now appear in multiple locations in the computational graph, multiplying their contribution without increasing storage requirements. During backpropagation, shared parameters recieve gradient signals from multiple positions, strengthening gradient flow.
Specifically, we choose a block inside a layer and share its parameters (along with their masks) with all blocks after it. These later blocks reference the single block tensor that they share with, and multiply by a learnable scaling factor for forward and backward passes. For example, in a ResNet layer with 4 blocks (Figure 4), if the second block shares parameters with blocks 3 and 4, a parameter in block 2 effectively appears three times in the computation graph while only being stored and optimized once.
Thus, we adjust sparsity calculation to account for weight sharing, such that , where is the -norm, is the unique learnable parameter count, and is the theoretical parameter count in the original network with a mask . Formally, let be the number of residual blocks in a layer, and be the block that shares parameters (), where the remaining blocks share parameters with the -th block. Let represent the parameters of the -th block. Our weight sharing scheme can be expressed as:
@@ 3.3 Cyclic Sparsity
While standard DST methods such as SET and RigL adjust the sparsity pattern, sparsity is consistent throughout training - pruning and regrowing of weights take place at the same . i.e., . In our work, we introduce a cyclic sparsity schedule where the network’s sparsity level itself evolves dynamically during training. Instead of maintaining a constant sparsity, we cyclically adjust the sparsity between a maximum value and a minimum value . The sparsity level is governed by a cosine schedule:
where is the period of the cyclic sparsity. This schedule ensures that the sparsity grows smoothly to and then gradually reduces to . As a result, the network experiences phases of higher and lower connectivity, promoting better parameter exploration during training. Once the cyclic phase ends at , we fix the sparsity and switch to a static pruning and growth update. The details of our method are summarized in Algorithm 1.
@@ 4 Experiments
Evaluation Protocol. Our experiments include image classification using ResNet-34 and ResNet-50 [12] on the benchmark datasets CIFAR-10 and CIFAR-100, as well as ResNet-50 on ImageNet, following the same evaluation protocol as other SOTA methods in terms of datasets and backbones. We compare our method with DST methods SET[25] and RigL[4], and a competitive SST method SynFlow[31], which also experiments at very high sparsities. For SynFlow, we use the official repository with its default hyperparameters to evaluate its performance. For SET and RigL, we use the repository provided by ITOP [21]. We keep all hyperparameters as they are optimally configured in the paper, including , ERK initialization, and the mask update interval of 1500 and 4000 for SET and RigL, respectively.
In addition, we benchmark DCT, the SOTA method specifically focused on extreme sparsity using the repository provided by the paper [28]. While this method performs very competitively at extreme sparsity settings, it comes at a cost due to the computation of a dense matrix matching the full architecture size, which adds an extra layer of computational overhead and requires extra memory. This method achieves the best performance when paired with RigL, which we evaluate, and then we combine it with our approach to evaluate if our method can improve it further. In this integration, we use only DyReLU phasing and cyclic density, leaving weight sharing to prevent conflicts with the DCT matrices embedded in the model architecture. Here, we use = 100 to match results from the original paper, and we also use this value for our integration to ensure fairness.
@@
| Synflow | 82.03 | 61.61 | 56.33 | 10.00 | 57.44 | 22.91 | 1.00 | 1.00 |
| SET | 93.09 | 82.70 | 70.84 | 10.00 | 71.97 | 47.64 | 32.14 | 1.00 |
| RigL | 92.92 | 85.71 | 81.47 | 10.03 | 71.72 | 50.26 | 40.92 | 1.14 |
| EAST (ours) | 93.03 | 86.76 | 83.83 | 70.57 | 72.01 | 52.16 | 42.53 | 18.84 |
| DCTpS+RigL | 89.04 | 83.96 | 80.80 | 77.38 | 63.12 | 51.30 | 45.68 | 37.59 |
| DCTpS+RigL+EAST (ours) | 89.22 | 84.19 | 81.96 | 77.54 | 64.33 | 53.21 | 47.80 | 38.02 |
@@
| Synflow | 77.06 | 48.62 | 10.00 | 10.00 | 48.91 | 12.07 | 9.40 | 1.00 |
| SET | 91.97 | 27.05 | 16.04 | 10.00 | 70.35 | 8.59 | 1.26 | 1.00 |
| RigL | 92.75 | 58.41 | 30.14 | 10.00 | 70.50 | 51.93 | 7.02 | 1.00 |
| EAST (ours) | 93.12 | 86.16 | 76.60 | 54.63 | 70.64 | 52.84 | 14.23 | 1.88 |
| DCTpS+RigL | 89.66 | 85.98 | 85.03 | 83.13 | 64.85 | 55.98 | 52.71 | 49.63 |
| DCTpS+RigL+EAST (ours) | 89.67 | 87.10 | 85.87 | 83.88 | 64.92 | 58.17 | 55.60 | 51.33 |
@@
CIFAR-10 and CIFAR-100. We evaluate EAST with ResNet-34 and ResNet-50 backbones, with a particular focus on extreme sparsities (99%-99.99%). We repeat our experiments three times and report the average accuracy in Table 1 (variance visualized in Figure 2), which presents results across different configurations.
At 99%, all dynamic training methods achieve comparable performance, significantly outperforming SynFlow. This is unsurprising as a static mask is expected to lead to worse performance than a dynamic mask [4, 26]. EAST’s advantages become increasingly pronounced at higher sparsities. At 99.90% and 99.95%, EAST beats the next best performing method RigL by a clear margin ranging between 1-2% with ResNet-34 on both datasets, showing significant gains in performance. The difference becomes even more clear with ResNet50, where other methods take a huge plummet on accuracy and EAST shows a more gradual decline. EAST displays an improvement of 38-40% over RigL on CIFAR-10, and 1-7% on CIFAR-100 at 99.90% and 99.95%. The most striking difference appears at 99.99%. We observe that all other methods result in complete collapse to random performance on either dataset and model. In contrast, EAST maintains meaningful accuracies except for ResNet50 on CIFAR-100. For instance, EAST achieves 70.57% on CIFAR-10 compared to 10% random baselines. This pattern holds across both models, though ResNet50 shows more degradation at the highest sparsity levels, presumably due to more prominent vanishing gradient problems in a deeper network, with even fewer parameters spread across each layer.
We evaluate DCTpS and find that while DCTpS demonstrates a clear advantage at 99.95% and 99.99%, it performs worse than SET, RigL and EAST at 99% and sometimes 99.9%. For instance, EAST achieves 3-8% higher accuracy than DCTpS at 99%, without involving additional overhead. To test the versatility of EAST, we further combine it with DCTpS and observe that EAST provides consistent improvements across all configurations tested.
@@
ImageNet. Prior methods like SynFlow and DCTpS were evaluated primarily on smaller datasets like CIFAR and Tiny-ImageNet. We extend these baselines to ImageNet by implementing and adapting their published code and evaluate the scalability of EAST on ImageNet using ResNet-50 at sparsity 99.5% and 99.9%. Given the much shortened training epochs for ImageNet, we implement only the DyReLU phasing and weight sharing components, excluding cyclic sparsity. As presented in Table 3, EAST has demonstrated capability in maintaining network robustness on large-scale dataset, outperforming previous SOTA by a significant margin, achieving approximately 2% and 2.7% higher top-1 accuracy than RigL. This advantage is more notable when comparing to other methods including SynFlow and SET, which fail to maintain meaningful performance and collapase to random accuracy.

### v4

@@ Abstract
Pruning of deep neural networks has been an effective technique for reducing model size while preserving most of the performance of dense networks, crucial for deploying models on memory and power-constrained devices. While recent sparse learning methods have shown promising performance up to moderate sparsity levels such as 95% and 98%, accuracy quickly deteriorates when pushing sparsities to extreme levels due to unique challenges such as fragile gradient flow. In this work, we explore network performance beyond the commonly studied sparsities, and develop techniques that encourage stable training without accuracy collapse even at extreme sparsities, including 99.90%, 99.95% and 99.99% on ResNet architectures. We propose three complementary techniques that enhance sparse training through different mechanisms: 1) Dynamic ReLU phasing, where DyReLU initially allows for richer parameter exploration before being gradually replaced by standard ReLU, 2) weight sharing which reuses parameters within a residual layer while maintaining the same number of learnable parameters, and 3) cyclic sparsity, where both sparsity levels and sparsity patterns evolve dynamically throughout training to better encourage parameter exploration. We evaluate our method, which we term Extreme Adaptive Sparse Training (EAST) at extreme sparsities using ResNet-34 and ResNet-50 on CIFAR-10, CIFAR-100, and ImageNet, achieving competitive or improved performance compared to existing methods, with notable gains at extreme sparsity levels. Code is available at https://github.com/TensorStrike/ES2.
@@ 1 Introduction
Network pruning (Han et al., 2015a; LeCun et al., 1990; Liu et al., 2017; Li et al., 2016; Kusupati et al., 2020; Han et al., 2015b) is a widely-used technique for reducing a network’s parameters and compressing its size. Reducing model sizes is crucial for deploying models on edge devices with limited resources. Conventionally, pruning methods have focused on reducing parameters from pre-trained models. However, it requires at least as much computation as training a dense model as it must converge before pruning takes place. The Lottery Ticket Hypothesis work (Frankle & Carbin, 2018; Frankle et al., 2020a; Malach et al., 2020) gives theoretical foundation that subnetworks have the potential to reach full performance even when trained from an initially sparse state. This insight has recently gained much traction in sparse training, a paradigm in which sparse networks are trained from scratch without the need for dense pre-training.
@@
Sparse training methods can be broadly classified into two categories: static sparse training (SST), also sometimes referred to as Pruning at Initialization (PaI), where the sparsity pattern is pre-determined at initialization and remains fixed throughout training (Lee et al., 2019; Wang et al., 2020; Tanaka et al., 2020a; De Jorge et al., 2020), and dynamic sparse training (DST), where the sparsity pattern continuously evolves during training (Mocanu et al., 2018; Dettmers & Zettlemoyer, 2019; Evci et al., 2021; Yin et al., 2022). These methods perform well at moderate sparsity commonly up to 98%, but experience rapid accuracy degradation beyond this due to limitations such as layer collapse and gradient issues (Tanaka et al., 2020b; Wang et al., 2020). This raises a fundamental question - what are the limits of sparse training? Understanding these boundaries is both scientifically valuable for characterizing sparse learning dynamics and potentially relevant for future deployment scenarios, where resource constraints continue to be problematic. However, extreme sparsities remain relatively understudied, with only a few methods attempting to maintain usable accuracy at these extreme levels (Tanaka et al., 2020b; De Jorge et al., 2020; Price & Tanner, 2021).
@@
Our work proposes a collection of adaptive methods that can achieve meaningful performance at extreme sparsities. By leveraging three core strategies - 1) phased Dynamic ReLU activation, 2) weight sharing, 3) cyclic sparsity scheduling - we present a flexible framework that is capable of optimizing performance at extreme sparsities. Each component independently enhances the sparse model’s capacity as seen in our ablation studies, where individual strategy improves performance on their own. These methods can also cohesively achieve an even stronger performance when combined. This framework, which we term Extreme Adaptive Sparse Training (EAST), offers a modular approach that can be adapted to existing DST frameworks, postponing their total performance collapse. Through empirical analysis, we find that EAST prevents gradient vanishing and encourages parameter exploration, achieving notable performance gain over our baselines. Our primary focus is on the final compressed model for deployment rather than training efficiency, as these represent distinct optimization objectives in resource-constrained scenarios. Our contribution and key achievements can be summarized as follows:
@@
We perform extensive empirical evaluation of EAST on image classification tasks, and demonstrate competitive performance at sparsities beyond 99.90%.
We explore and demonstrate that DyReLU when used as a drop-in replacement for ReLU at initialization and then replaced completely can still improve the parameter exploration at extreme sparsities, leading to higher performance than sparse models trained using ReLU only.
Unlike existing DST methods that maintain fixed sparsity levels, EAST introduces cyclic sparsity scheduling where both patterns and sparsity levels evolve dynamically during training. Our cyclic sparsity schedule allows models to explore parameters and stabilize at extreme sparsity, as evidenced by our ablation studies.
@@ 2 Related Work
Dynamic Sparse Training. Early work by Mocanu et al. (2018) demonstrated the viability of DST through SET, which used magnitude pruning and random weight regrowth. Subsequent criteria for regrowing connections have been developed, such as using the momentum of parameters in SNFS (Dettmers & Zettlemoyer, 2019) or their gradient in RigL (Evci et al., 2021) to determine the salience of deactivated connections. GraNet (Liu et al., 2021a) adapted RigL’s approach by starting with half the parameters and gradually increasing to target sparsity instead of starting from the target sparsity. ITOP (Liu et al., 2021b) provides a deeper understanding of pre-existing DST methods and enhances parameter exploration during sparse training. Supticket (Yin et al., 2022) further enhances the performance of DST by leveraging weight averaging at the late training stage. NeurRev (Li et al., 2024) addressed the issue of dormant neurons by removing harmful negative weights. Recent advances include Top-KAST (Jayakumar et al., 2020), which maintains an auxiliary set of weights for enhanced exploration, and BiDST (Ji et al., 2024) which formulates DST as a bi-level optimization problem for joint weight-mask optimization. MEST (Yuan et al., 2021) introduces a memory-efficient sparse training framework for edge devices, while Chase (Yin et al., 2024) adapts dynamic unstructured sparsity into hardware-friendly channel-wise sparsity, both optimizing performance on resource-constrained systems without sacrificing accuracy. SRigL Lasby et al. (2024) extends the RigL-based pipeline to structured sparsity and imposes a constant fan-in N:M for better hardware efficiency. CHT Zhang et al. (2025) adopts a regrowth strategy that does not require the use of gradient information in backpropagation, and solely relies on topological information instead, allowing the regrowth computation to be cheaper. DSCR Wu et al. (2025) shows that, at moderate sparsity, dynamically sparse models can outperform dense ones in robustness to common image corruptions. DynaDiag Wu et al. (2025) restricts parameters to a diagonal sparsity pattern, which preserves full input–output coverage while enabling GPU-friendly structured sparsity and dynamic mask updates within that pattern.
@@
Extreme Sparsity Methods. Most current methods focus on sparsity levels at which pruned networks match or closely approximate the performance of their dense counterparts, a concept referred to as matching sparsity (Frankle et al., 2020b). However, there is growing interest in extreme sparsities, where accuracy trade-offs become inevitable but potentially worthwhile for substantial computational gains. Yet there remain challenges, such as maintaining stable gradient flow and preventing layer collapse, which conventional pruning methods struggle to address. GraSP (Wang et al., 2020) and SynFlow (Tanaka et al., 2020b) specifically addressed the gradient preservation and layer collapse issues at high sparsities via careful network initialization. SNIP (Lee et al., 2019) and SNIP-it (Verdenius et al., 2020) proposed PaI methods that showed promising results at higher sparsities, though primarily focused on static masks. De Jorge et al. (2020) demonstrated that methods like SNIP and GraSP can perform worse than random pruning at 99% sparsity and beyond, and proposed iterative pruning approaches (Iter-SNIP and FORCE) that maintained meaningful performance up to 99.5% sparsity through gradual parameter elimination. DCTpS (Price & Tanner, 2021) introduced "DCT plus Sparse" layers that combined a fixed dense DCT offset matrix with trainable sparse parameters to maintain propagation at extreme sparsities up to 99.99%.
@@
Activation Functions. Activation functions in sparse networks have traditionally relied on static activation functions such as ReLU (Nair & Hinton, 2010), which applies a fixed threshold to the input. Parametric variants such as PReLU (He et al., 2015) introduced learnable parameters to adapt the negative slope, offering slight improvements in gradient flow. Swish (Ramachandran et al., 2017) and Mish (Misra, 2019) aim to smooth gradients for better optimization but remain static throughout training. Dynamic ReLU (DyReLU) (Chen et al., 2020) introduced an adaptive activation that dynamically adjusts the slopes of a ReLU based on the input.
@@ 3.1 Activation Function
In this work, when initializing a network, we replace the standard ReLU activation with a non-ReLU activation. We tested several activations and found that DyReLU performed best, specifically the DyReLU-B variant, which dynamically adapts and adjusts channel-wise activation coefficients during training, enabling richer expressivity during early training. As training progresses, we gradually transition from DyReLU to standard ReLU, eliminating the additional parameters. Preliminary details about DST, DyReLU and details on phasing out DyReLU can be found in the Supplementary Materials.
Understanding Gradient Flow Through Neuron Activity. With ReLU, non-positive pre-activations would result in zero gradient and prevent the learning signal from propagating. To better understand why DyReLU phasing improves performance at extreme sparsity, we analyzed pre-activation values throughout training. As shown in Figure 3, the model trained solely with ReLU maintains around 1.3% positive pre-activations throughout training, with many deeper layers showing none, interrupting gradient propagation.
@@
This mechanism of preserving activation pathways directly translates to sustained gradient flow during training. As demonstrated in Figure 5, networks trained with DyReLU phasing maintain healthy gradient norms even after complete conversion to standard ReLU, whereas networks trained solely with ReLU exhibit near-zero gradient flow and complete performance collapse.
Phasing Out DyReLU. As training progresses, we phase out all DyReLU and gradually replace them with the standard ReLU. We introduce a hyperparameter that regulates the contribution of DyReLU during training. It makes full contribution when it is 1, and gradually decreases to 0 as training progresses, eventually reaching 0 before the first learning rate decay, at which point DyReLU has been completely replaced by ReLU. We find that this helps the network settle into a more stable configuration, as DyReLU is no longer needed once the network has explored the parameter space adequately. This process removes all DyReLU-introduced parameters.
By the end of phasing , we have:
@@ 3.2 Weight Sharing
Much research in sparse training has been about mask optimization - determining the best subset of parameters to keep. However, at extreme sparsities, gradient flow can collapse even with an optimal parameter subset. Therefore, we ask a different question: Can we increase the number of parameter paths in the computational graph while maintaining the same number of learnable parameters? Our approach creates multiple references to the same parameter tensors, effectively allowing one learnable parameter to contribute to multiple locations in the forward/backward pass without additional storage cost.
Our weight sharing mechanism strategically reuses parameters within each layer’s residual blocks, allowing the remaining active parameters to participate in multiple paths of the computational graph. If, at a certain sparsity, we have remaining parameters in the original model, we keep exactly this many unique parameters to learn, but some of them now appear in multiple locations in the computational graph, multiplying their contributions without increasing storage requirements. During backpropagation, gradients from multiple blocks sharing the same parameters are automatically summed by PyTorch’s autograd, strenghening the learning signal for these weights. Each sharing block includes learnable scaling factors that allow the blocks to specialize while maintaining parameter efficiency.
Specifically, we choose a block within a layer and share its parameters (along with their masks) with all blocks that follow it. These later blocks reference the single block tensor that they share with, and multiply by a learnable scaling factor for forward and backward passes. For example, in a ResNet layer with 4 blocks (Figure 4), if the second block shares parameters with block 3 and 4, a parameter in block 2 effectively appears three times in the computation graph while only being stored and optimized once.
Since multiple computational paths now reference the same learnable parameters, we adjust our sparsity calculation to reflect the true number of unique learnable parameters rather than the apparent network size. Thus, , where represents only the unique learnable parameter count being optimized. is the theoretical parameter count in the original network with a mask . Formally, let be the number of residual blocks in a layer, and be the block that shares parameters (), where the remaining blocks share parameters with the -th block. Let represent the parameters of the -th block. Our weight sharing scheme can be expressed as:
@@ 3.3 Cyclic Sparsity
While classic DST methods such as SET and RigL adjust the sparsity pattern, sparsity remains constant throughout training - pruning and regrowing of weights take place at the same rate. i.e., . In our work, we introduce a cyclic sparsity schedule where the network’s sparsity level itself evolves dynamically during training. Instead of maintaining constant sparsity, we cyclically adjust the sparsity between a maximum value and a minimum value over a period . This creates phases of higher and lower connectivity, allowing the network to explore parameters more effectively, preventing the gradient issues that plague static extreme sparsity training. Once the cyclic phase ends at , we fix the sparsity at and switch to standard DST updates. Details of the method are summarized in Algorithm 1.
@@ 4 Experiments
Evaluation Protocol. Our experiments include image classification using ResNet-34 and ResNet-50 (He et al., 2016) on the benchmark datasets CIFAR-10 and CIFAR-100, as well as ResNet-50 on ImageNet, following the same evaluation protocol as other SOTA methods in terms of datasets and backbones. We compare our method with DST methods SET (Mocanu et al., 2018) and RigL (Evci et al., 2021), and a competitive SST method SynFlow (Tanaka et al., 2020b), which also experiments at very high sparsities. For SynFlow, we use the official repository with its default hyperparameters to evaluate its performance. For SET and RigL, we use the repository provided by ITOP (Liu et al., 2021b). We keep all hyperparameters as they are optimally configured in the paper, including , ERK initialization, and the mask update interval of 1500 and 4000 for SET and RigL, respectively.
In addition, we benchmark DCT, the SOTA method specifically focused on extreme sparsity, using the repository provided by the method in Price & Tanner (2021). While this method performs very competitively at extreme sparsity settings, it comes at a cost due to the computation of a dense matrix matching the full architecture size, which adds an extra layer of computational overhead and requires extra memory. This method achieves the best performance when paired with RigL, which we evaluate, and then we combine it with our approach to evaluate if our method can improve it further. In this integration, we use only DyReLU phasing and cyclic density, leaving weight sharing to prevent conflicts with the DCT matrices embedded in the model architecture. Here, we use = 100 to match results from the original paper, and we also use this value for our integration to ensure fairness.
@@
| Synflow | 86.030.71 | 61.6110.76 | 56.333.44 | 10.000.00 | 57.441.72 | 22.9118.98 | 1.000.00 | 1.000.00 |
| SET | 93.090.15 | 82.700.91 | 70.841.51 | 10.000.00 | 71.970.72 | 47.640.43 | 32.140.60 | 1.000.00 |
| RigL | 92.920.18 | 85.710.23 | 81.470.32 | 10.030.19 | 70.720.25 | 50.260.16 | 38.440.95 | 1.140.24 |
| EAST (ours) | 93.510.13 | 86.990.32 | 83.830.02 | 62.120.90 | 71.140.51 | 56.32 0.31 | 45.810.60 | 18.840.55 |
| DCTpS+RigL | 89.040.07 | 83.960.23 | 80.800.51 | 77.380.57 | 63.120.53 | 51.300.26 | 45.680.70 | 37.590.52 |
| DCTpS+RigL+EAST (ours) | 89.720.09 | 84.190.64 | 81.960.24 | 77.540.17 | 64.330.44 | 53.210.28 | 47.800.69 | 38.020.26 |
@@
| Synflow | 77.067.40 | 48.629.32 | 10.000.00 | 10.000.00 | 48.911.48 | 12.0710.21 | 9.403.92 | 1.000.00 |
| SET | 91.970.25 | 27.0515.17 | 16.042.85 | 10.000.00 | 70.350.11 | 8.592.61 | 1.260.28 | 1.000.00 |
| RigL | 92.750.25 | 58.4142.03 | 30.1410.85 | 10.000.00 | 70.500.32 | 51.900.62 | 7.025.33 | 1.000.00 |
| EAST (ours) | 93.480.13 | 86.160.12 | 81.180.59 | 49.554.69 | 71.040.15 | 56.760.22 | 45.550.43 | 3.642.84 |
| DCTpS+RigL | 89.660.27 | 85.980.55 | 85.030.25 | 83.130.53 | 64.850.23 | 55.980.09 | 52.710.20 | 49.631.11 |
| DCTpS+RigL+EAST (ours) | 89.670.12 | 87.100.35 | 85.870.39 | 83.880.19 | 64.921.41 | 58.170.22 | 55.600.38 | 51.330.57 |
@@
CIFAR-10 and CIFAR-100. We evaluate EAST with ResNet-34 and ResNet-50 backbones, with a particular focus on extreme sparsities (99%-99.99%). We repeat our experiments three times and report the average accuracy in Table 1, which presents results across different configurations.
At 99%, all dynamic training methods achieve comparable performance, significantly outperforming SynFlow. This is unsurprising as a static mask is expected to lead to worse performance than a dynamic mask (Evci et al., 2021; Mostafa & Wang, 2019). EAST’s advantages become increasingly pronounced at higher sparsities.
At 99.90% and 99.95%, EAST consistently outperforms existing methods across both architectures and datasets, with particularly notable improvements on ResNet-50 where competing methods experience severe accuracy degradation. While other approaches show dramatic performance collapse beyond 99.90% sparsity, EAST demonstrates more gradual degradation patterns.
At 99.99%, we observe that all other methods result in complete collapse to random performance across both datasets and architectures. In contrast, EAST maintains non-random accuracies. ResNet-50 shows greater sensitivity at this extreme sparsity, likely due to vanishing gradient issues exacerbated by the deeper architecture and reduced parameter density per layer.
Our evaluation of DCTpS reveals complementary strengths. While DCTpS excels at the highest sparsity levels (99.95% and 99.99%), classic DST methods including EAST maintain advantages at moderate extreme sparsities (99% and 99.9%) without requiring additional computational overhead. When combined with DCTpS, EAST consistently improves across all tested configurations, demonstrating the modular nature of our approach and its ability to enhance existing extreme sparsity methods.
ImageNet. Prior methods, such as SynFlow and DCTpS, were primarily evaluated on smaller datasets, such as CIFAR and Tiny-ImageNet. We extend these baselines to ImageNet by implementing and adapting their published code and evaluate the scalability of EAST on ImageNet using ResNet-50 at sparsity 99.5% and 99.9%. Given the much shortened training epochs for ImageNet, we implement only the DyReLU phasing and weight sharing components, excluding cyclic sparsity. As presented in Table 3, EAST has demonstrated scalability to large-scale datasets compared to existing methods. While SynFlow and SET collapse to random accuracies, EAST maintains stable performance across both tested sparsity levels and outperforms RigL by a notable margin.